# Adaptively Bypassing Vision Transformer Blocks for Efficient Visual Tracking

### v1

@@ Abstract
Empowered by transformer-based models, visual tracking has advanced significantly. However, the slow speed of current trackers limits their applicability on devices with constrained computational resources. To address this challenge, we introduce ABTrack, an adaptive computation framework that adaptively bypassing transformer blocks for efficient visual tracking. The rationale behind ABTrack is rooted in the observation that semantic features or relations do not uniformly impact the tracking task across all abstraction levels. Instead, this impact varies based on the characteristics of the target and the scene it occupies. Consequently, disregarding insignificant semantic features or relations at certain abstraction levels may not significantly affect the tracking accuracy. We propose a Bypass Decision Module (BDM) to determine if a transformer block should be bypassed, which adaptively simplifies the architecture of ViTs and thus speeds up the inference process. To counteract the time cost incurred by the BDMs and further enhance the efficiency of ViTs, we innovatively adapt a pruning technique to reduce the dimension of the latent representation of tokens in each transformer block. Extensive experiments on multiple tracking benchmarks validate the effectiveness and generality of the proposed method and show that it achieves state-of-the-art performance. Code is released at: https://github.com/1HykhqV3rU/ABTrack.
Keywords Efficient Visual Tracking, Adaptively Bypassing, Pruning
Figure 1: The attention mechanism in ViTs transforms image patch representation vectors repeatedly, progressively incorporating semantic relationships between image patches through a series of transformer blocks. However, deep semantic features or relations are not always necessary for visual tracking task. This figure shows the tracking results of three trackers adapted from DropMAE with different number of ViT layers, dubbed DropMAE-$N$, $N=1,6,11$ being the number of layers. As shown, despite DropMAE-11 successfully tracking all three targets, DropMAE-1 succeeds in tracking the boat with only one layer and DropMAE-6 succeeds in tracking the boat and the person with only six layers. These results suggest that low-level semantic features or relations are adequate for certain tracking tasks. This observation inspires our approach of adaptively bypassing ViT blocks for improving tracking efficiency.
@@ 1 Introduction
Visual object tracking is a foundational task in computer vision that entails monitoring an object’s movement in a video sequence from its initial state Li et al. (2018a); Jiao et al. (2021); Javed et al. (2021). Recent advancements in deep neural networks, including AlexNet Krizhevsky et al. (2012), ResNet He et al. (2016), GoogLeNet Szegedy et al. (2015), and transformers Vaswani et al. (2017), have significantly enhanced tracking performance. Transformers, in particular, have been instrumental in developing high-performance trackers  Chen et al. (2021a); Yan et al. (2021a); Wang et al. (2021a); Yu et al. (2021); Cui et al. (2022a); Ye et al. (2022a); Chen et al. (2023). However, many existing tracking methods have primarily focused on achieving top-notch performance on powerful GPUs, often neglecting considerations of tracking speed Li et al. (2018b); Bhat et al. (2019); Chen et al. (2021a). While many of these methods may deliver real-time performance on robust hardware, they frequently fall short in terms of competitiveness and adaptability to resource-constrained devices.
Recently, the tracking community has observed a surge in the adoption of single-stream architectures, where feature extraction and fusion are seamlessly integrated via harnessing the capabilities of pre-trained ViT backbone networks Ye et al. (2022a); Chen et al. (2022); Xie et al. (2022); Cui et al. (2022a). In our study, we also employ a single-stream architecture, but our focus lies in enhancing the efficiency of ViTs for more effective visual tracking. As is known, the attention mechanism in ViTs transforms image patch representation vectors repeatedly, progressively incorporating semantic relationships between image patches through a series of transformer blocks Dosovitskiy et al. (2021); Liu et al. (2021); Khan et al. (2021). This capability empowers ViTs to extensively capture both local features or relations within objects and broader-scale ones between objects and objects/background, thereby significantly enhancing their performance in image classification tasks Wang et al. (2021b); Arnab et al. (2021); Mehta and Rastegari (2022); Han et al. (2020). Nevertheless, a notable disparity exists between visual tracking and image classification.
@@
In image classification, local or broader-scale features or relations crucial for capturing class characteristics or specific qualities are essentially intrinsic to the object, unaffected by surrounding circumstances. However, in visual tracking, the significance of semantic features or relations exhibits a dynamic nature. The variation in impact is intricately tied to the specific characteristics of both the target and the surrounding environment. In other words, the tracking process is not solely dictated by a fixed set of semantic features or relations across all levels of abstraction. Instead, it operates as a dynamic and context-sensitive mechanism, involving continuous adaptation and interpretation of dynamic changes in the object’s surroundings. This dynamic nature of tracking necessitates a flexible and context-sensitive approach, where the tracker must dynamically adjust its focus and attention based on the evolving context of the scene. For instance, in scenarios where a target moves against a monochromatic background, simply the color contrast between the target and the background can be sufficient for effective tracking. However, when the target is moving in a cluttered environment, capturing adequate semantic features and relations becomes crucial for achieving effective tracking (see Figure 1 for an intuitive illustration). This observation underscores the need for a tracking methodology that adapts to the intricacies of each tracking scenario, recognizing that the significance of semantic features or relations is context-dependent. Unfortunately, this insightful perspective has been largely overlooked in previous works.
@@
In this work, we propose an adaptive computation framework for efficient visual tracking based on the above observation. Given that the difficulty of tracking a target depends on the background it occupies, we propose to adaptively bypass certain transformer blocks. This allows for the selective extraction of significant semantic features or relations tailored to specific tracking tasks, while disregarding blocks representing insignificant, negligible or repeated features or relations. To achieve this, we introduce a bypass decision module into each transformer block. This module outputs a bypass probability, which determines whether a transformer block should be bypassed. In our implementation, the bypass decision module is a linear layer followed by a nonlinear activation function, which takes an additionally introduced bypass token as input. By tailoring the architecture of ViTs to the specific demands of a tracking task, our approach holds the potential to accelerate the inference process for visual tracking. To counteract the computational burden incurred by the these introduced modules and further improve the efficiency of ViTs, we also reduce the dimension of latent representation of tokens by introducing a novel adaptation of an existing ViT pruning method proposed in Zhu et al. (2021). This adjustment includes a local ranking of importance scores and an extended applicability due to our generalized formulation. We called this proposed adaptive computation framework ABTrack. Extensive experiments substantiate the effectiveness and generality of our method and demonstrate that our ABTrack achieves state-of-the-art real-time performance. Our contributions can be summarized as follows:
@@
  • To counteract the time cost of BDM and further enhance efficiency, we propose to reduce the dimension of latent representations of tokens by innovatively adapting an existing ViT pruning method.
@@ 2.1 Visual Tracking
Siamese-based approaches, such as SiamFC Bertinetto et al. (2016), SiamRPN Li et al. (2018b), SiamCAR Guo et al. (2020), and SiamBAN Chen et al. (2020), are widely used in visual tracking. This framework typically utilizes two backbone networks sharing parameters to extract features from the template and search region images. It employs a correlation-based network to interact with the features and uses head networks for prediction. To further enhance tracking performance, TransT Chen et al. (2021a), TrSiam Wang et al. (2021a), and their subsequent works such as Stark Yan et al. (2021a), ToMP Mayer et al. (2022), CSWinTT Song et al. (2022), AiATrack Gao et al. (2022), and GRM Gao et al. (2023) introduce the transformer for feature interaction. Recently, a single-stream framework, exemplified by such as MixFormer Cui et al. (2022a), OSTrack Ye et al. (2022a), SimTrack Chen et al. (2022), DropTrack Wu et al. (2023), CTTrack Song et al. (2023), and ARTrack Wei et al. (2023a), has demonstrated great success in visual tracking. This framework combines feature extraction and fusion within a single backbone network. It is a simple yet effective approach that capitalizes on the capabilities of pre-trained image classification models. However, these methods were primarily designed for powerful GPUs, resulting in slow speeds on edge devices, which limits their practical applications. In our work, we also adopt this single-stream framework but with a focus on improving its efficiency.
@@ 2.2 Efficient Tracking Methods
Practical applications necessitate the use of efficient tracking systems capable of delivering exceptional performance at rapid speeds, especially when deployed on edge devices. While early techniques like ECO Danelljan et al. (2017) and ATOM Danelljan et al. (2019) were able to achieve real-time tracking on edge devices, their performance fell short in comparison to state-of-the-art trackers available today. Recent advancements have led to the development of more efficient tracking solutions. LightTrack Yan et al. (2021b), for instance, employs neural architecture search (NAS) to discover network configurations, resulting in a low computational overhead and relatively high performance. Afterwards, FEAR Borsuk et al. (2022) introduces a family of efficient and accurate tracking methods through the utilization of a dual-template representation and a pixel-wise fusion block. Recently, HiT Kang et al. (2023) presents a family of efficient transformer-based tracking models that use a bridge module and dual-image position encoding to bridge the gap between lightweight transformers and the tracking framework. MixFormerV2  Cui et al. (2023) proposes a fully transformer-based tracking framework without any dense convolutional operations or complex score prediction modules, additionally, it employs a distillation-based model compression paradigm to enhance its efficiency. Despite these advancements, these methods are tailored and have problems with universality. A widely applicable method that can be easily integrated into existing trackers is more desirable. In this work, we present a novel approach that is straightforward and can be seamlessly integrated into existing ViT-based methods.
@@ 2.3 Efficient Vision Transformers
ViT Dosovitskiy et al. (2021) revolutionized the field of computer vision with its impressive performance Touvron et al. (2021); Yuan et al. (2021); Wu et al. (2021); Wang et al. (2021b); Liu et al. (2021). However, despite their powerful modeling capabilities, they have been facing challenges associated with speed limitations. In situations prioritizing fast inference, especially on resource-constrained edge devices, the demand for efficient ViTs intensifies. To expedite ViTs, various lightweight ViTs have been recently introduced, employing techniques like low-rank methods, model compression, and hybrid designs Wang et al. (2020); Zhang et al. (2022); Mao et al. (2021); Chen et al. (2021b); Li et al. (2022). However, low-rank and quantization-based ViTs often sacrifice considerable accuracy for efficiency, while Hybrid ViTs with CNN-based components exhibit restricted input flexibility. In this work, we exploit pruning to reduce the dimension of latent representation of tokens, by innovatively adapting the ViT pruning method proposed in  Zhu et al. (2021). On the other hand, recent endeavors in efficient ViTs based on conditional computation have explored adaptive inference methods for accelerated modeling. For instance, DynamicViT Rao et al. (2021) introduces extra control gates to pause specific tokens using the Gumbel-softmax trick. A-ViT Yin et al. (2022) uses an Adaptive Computation Time (ACT)-like method to enhance efficiency, accuracy, and token importance allocation without requiring additional halting sub-networks. TPS Wei et al. (2023b) proposes a joint Token Pruning and Squeezing module for compressing vision transformers with higher efficiency. However, variable number of tokens in these methods incurs considerable time cost due to additional unstructured assess operations. In this study, we investigate adaptively bypassing ViT blocks as a structured and more effective method of conditional computation to enhance ViTs’ efficiency. As the tracking challenge for a specific target is highly context-dependent, adaptive bypassing of ViT blocks can notably decrease insignificant semantic features or relations at specific abstraction levels without significantly impacting the tracker’s performance.
Figure 2: Overview of the proposed ABTrack framework. It contains a single-stream backbone and a prediction head, in which the backbone consists of pruned ViT blocks and Bypass Decision Modules.
3 Method
In this section, we first briefly overview our tracking framework ABTrack, as illustrated in Figure 2. Then, we introduce the bypass decision module and the method of pruning vision transformer models. Finally, we detail the prediction head and training objective.
@@ 3.1 Overview
The proposed ABTrack is a single-stream tracking framework, consisting of an adaptive ViT-based backbone and a prediction head. The inputs to the framework are the target template and the search image. The extracted features from the backbone are fed into the prediction head to obtain tracking predictions. In the backbone network, we introduce a bypass decision module (BDM) for each ViT block (layer) except for the first $n_{enf}$ blocks, which is trained to adaptively determines whether to bypass the associated block based on the computed bypassing probability. Furthermore, we conduct pruning on the ViT blocks using two introduced dimension reduction (DR) matrices. These matrices indicate the importance scores for each dimension of the linear layer of the Multi-Head Self-Attention (MSA) and Multi-Layer Perceptron (MLP) within each block, respectively. Since the Bypass Decision Module (BDM) does not change the transform characteristics of each ViT layer, while pruning alters the input domain of the BDM and consequently impacts its performance, we conduct ViT pruning before integrating the BDMs. Finally, with BDMs integrated, the pruned model undergoes fine-tuning within the tracking framework.
@@ 3.2 Bypass Decision Module (BDM)
The objective of the BDM is to dynamically determine whether a transformer block should be bypassed based on the given template and search image. In consideration of efficiency, it is implemented by a linear layer followed by a nonlinear activation function. Its input is a bypass token, a specific summary token that is appended to the sequence of image tokens. And its output represents the probability of bypassing the associated transformer block. Specifically, take the $i_{th}$ ($i>n_{enf}$) layer for example, let the bypass token be denoted by $(t^{i}_{1},...,t^{i}_{d}):=\textbf{b}^{i}\in\mathbb{R}^{d}$, the linear layer be denoted by $\mathfrak{l}^{i}$, where $d$ is the token dimension, then formally the BDM at layer $i$ is defined by
$p_{i}=\mathfrak{s}(\mathfrak{l}^{i}(\textbf{b}^{i})),$ (1)
@@
where $p_{i}\in[0,1]$ represents the probability of bypassing the $i$ transformer block, $\mathfrak{s}(x)=1/(1+e^{-x})$ denotes the sigmoid function. If $p_{i}>\rho$, where $\rho\in[0,1]$ is the bypass probability threshold, the transformer block at layer $i$ is bypassed; otherwise, the output tokens from the $(i-1)_{th}$ layer is fed into the $i_{th}$ transformer block. Let $N$ denote the total number of transformer blocks of the given ViT. Theoretically, all $N$ blocks may be bypassed simultaneously, in which case no correlation between template and search image is computed. To avoid such ill conditions, the first $n_{enf}$ layers are enforced without bypassing, which may reduce computing burdens due to BDM since these low level blocks are presumably expected in most cases by virtue of that low level features provide foundational information that more complex features and representations can be built upon. An other extreme case is that, for whatever input, few or no blocks are bypassed as by which the model has more powerful representation ability and thus is easier to reduce the training loss. To address this issue, we propose an adaptive block sparsity loss, denoted as $\mathcal{L}_{spar}$. This loss function adaptively penalizes smaller average probabilities across all considered layers, with a stricter penalty imposed when the tracking task is easy and a less strict penalty when the task is challenging. This adaptive approach ensures that blocks are bypassed dynamically, striking a balance between efficiency and accuracy based on the complexity of the tracking scenario. The block sparsity loss is formally defined by
@@
$\begin{split}&\mathcal{L}_{spar}=|\frac{1}{N-n_{enf}}\sum^{N}_{i=n_{enf}+1}p_{i}-\tau|,\\ &\tau=clip(\tau_{0}+\varrho(\mathcal{L}_{iou}-\overline{\mathcal{L}_{iou}}),0,1),\end{split}$ (2)
@@
where $\text{clip}(x,a,b)$ is a mathematical operation that limits the value of $x$ within a specified range defined by $a$ and $b$. $\tau\in[0,1]$ is a function of GIoU loss Rezatofighi et al. (2019) (see Section 3.4) and is used along with $\rho$ to control the block sparsity. Note that a linear function of the deviation of GIoU loss from its average is used to quantify the difficulty of the tracking task. In our implementation, the average GIoU loss, i.e., $\overline{\mathcal{L}_{iou}}$, is estimated with a batch training samples. $\varrho>0$ and $\tau_{0}\in[0,1]$ are two constants used to model the linear relation between $\tau$ and $(\mathcal{L}_{iou}-\overline{\mathcal{L}_{iou}})$. We call $\tau$ the block sparsity function. Given $\rho$, generally the larger $\tau$ is, the more sparse the model will be.

### v2

@@ Abstract
Empowered by transformer-based models, visual tracking has advanced significantly. However, the slow speed of current trackers limits their applicability on devices with constrained computational resources. To address this challenge, we introduce ABTrack, an adaptive computation framework that adaptively bypassing transformer blocks for efficient visual tracking. The rationale behind ABTrack is rooted in the observation that semantic features or relations do not uniformly impact the tracking task across all abstraction levels. Instead, this impact varies based on the characteristics of the target and the scene it occupies. Consequently, disregarding insignificant semantic features or relations at certain abstraction levels may not significantly affect the tracking accuracy. We propose a Bypass Decision Module (BDM) to determine if a transformer block should be bypassed, which adaptively simplifies the architecture of ViTs and thus speeds up the inference process. To counteract the time cost incurred by the BDMs and further enhance the efficiency of ViTs, we introduce a novel ViT pruning method to reduce the dimension of the latent representation of tokens in each transformer block. Extensive experiments on multiple tracking benchmarks validate the effectiveness and generality of the proposed method and show that it achieves state-of-the-art performance. Code is released at: https://github.com/xyyang317/ABTrack.
keywords:
Efficient Visual Tracking, Adaptively Bypassing, Pruning
journal: Pattern Recognition
[…]
@@ 1 Introduction
Visual object tracking is a foundational task in computer vision, involving the continuous monitoring of an object’s movement within a video sequence from its initial state [1]. Over the past decade, significant advancements in deep neural networks have propelled this field forward. Networks such as AlexNet [2], ResNet [3], GoogLeNet [4], alongside the recent integration of transformers [5], have collectively enhanced the accuracy and robustness of tracking systems. Transformers, in particular, have been pivotal in the development of high-performance trackers [6, 7, 8, 9, 10], offering substantial improvements in tracking precision and reliability. Despite these technological leaps, there is a noticeable gap in the balance between performance and efficiency. Many state-of-the-art tracking methods are optimized for top-tier performance on powerful GPUs, often at the expense of tracking speed [6, 11, 12, 13]. These methods, while achieving real-time performance on high-end hardware, frequently struggle to maintain the same level of efficiency on resource-constrained devices. This highlights the need for further research to develop solutions that are both accurate and efficient, making them adaptable to various hardware capabilities.
Figure 1: This figure shows the tracking results of three trackers adapted from DropMAE [23] with different number of ViT layers, dubbed DropMAE-$N$, $N=1,6,11$ being the number of layers. Note that despite DropMAE-11 successfully tracking all three targets, DropMAE-1 succeeds in tracking the boat with only one layer and DropMAE-6 succeeds in tracking the boat and the person with only six layers. These results suggest that deep semantic features or relations are not always necessary for visual tracking task.
@@
Recently, the tracking community has observed a surge in the adoption of single-stream architectures, where feature extraction and fusion are seamlessly integrated via harnessing the capabilities of pre-trained ViT backbone networks [10, 14, 9]. In our study, we also employ a single-stream architecture, but our focus lies in enhancing the efficiency of ViTs for more effective visual tracking. The attention mechanism in ViTs transforms image patch representation vectors repeatedly, progressively incorporating semantic relationships between image patches through a series of transformer blocks [15, 16]. This capability empowers ViTs to extensively capture both local features or relations within objects and broader-scale ones between objects and objects/background, thereby significantly enhancing their performance in image classification tasks [17, 18, 19]. However, a critical distinction exists between visual tracking and image classification, which necessitates further investigation. In image classification, the local and broader-scale features or relationships essential for capturing class characteristics are intrinsic to the object and remain unaffected by surrounding circumstances. In contrast, visual tracking requires a dynamic understanding of semantic features and relationships. The significance of these features is intricately tied to the specific characteristics of both the target and the surrounding environment. In other words, the tracking process is not solely dictated by a fixed set of semantic features or relations across all levels of abstraction. Instead, it operates as a dynamic and context-sensitive mechanism, involving continuous adaptation and interpretation of dynamic changes in the object’s surroundings. This dynamic nature of tracking necessitates a flexible and context-sensitive approach, where the tracker must dynamically adjust its focus and attention based on the evolving context of the scene. For instance, in scenarios where a target moves against a monochromatic background, simply the color contrast between the target and the background can be sufficient for effective tracking. However, when the target is moving in a cluttered environment, capturing adequate semantic features and relations becomes crucial for achieving effective tracking (see Figure 1 for an intuitive illustration). This observation underscores the need for a tracking methodology that adapts to the intricacies of each tracking scenario, recognizing that the significance of semantic features or relations is context-dependent. Unfortunately, this insightful perspective has been largely neglected in visual tracking research. More comprehensive explorations could be both fruitful and worthwhile, potentially leading to significant advancements in the field.
@@
In this work, we propose an adaptive computation framework for efficient visual tracking based on the above observation. Given that the difficulty of tracking a target depends on the background it occupies, we propose to adaptively bypass certain transformer blocks. This allows for the selective extraction of significant semantic features or relations tailored to specific tracking tasks, while disregarding blocks representing insignificant, negligible or repeated features or relations. To achieve this, we introduce a bypass decision module into each transformer block. This module outputs a bypass probability, which determines whether a transformer block should be bypassed. In our implementation, the bypass decision module is a linear layer followed by a nonlinear activation function, which takes an introduced bypass token as input. By tailoring the architecture of ViTs to the specific demands of a tracking task, our approach holds the potential to accelerate the inference process for visual tracking. To counteract the computational burden incurred by these introduced modules and further improve the efficiency of ViTs, we also reduce the dimension of latent representation of tokens by introducing a novel ViT pruning method. This innovative adjustment includes a local ranking of importance scores and an extended applicability due to our generalized formulation. We called this proposed adaptive computation framework ABTrack. Extensive experiments substantiate the effectiveness and generality of our method and demonstrate that our ABTrack achieves state-of-the-art real-time performance. Our contributions can be summarized as follows:
  • 1.
@@
  • 2.
    To counteract the time cost of BDM and further enhance efficiency, we propose to reduce the dimension of latent representations of tokens by introducing a novel ViT pruning method based on local ranking.
  • 3.
@@
The rest of this paper is organized as follows. Section 2 reviews related research. Section 3 details the methodology of the proposed technique. Section 4 describes the experiments conducted and discusses the results. Lastly, Section 6 presents the conclusions and insights drawn from the study.
@@ 2.1 Visual Tracking
Siamese-based approaches, such as SiamFC [21] and SiamRPN [11] are widely used in visual tracking. This framework typically utilizes two backbone networks sharing parameters to extract features from the template and search region images. It employs a correlation-based network to interact with the features and uses head networks for prediction. To further enhance tracking performance, TransT [6], TrSiam [8], and their subsequent works such as Stark [7], ToMP [12], CSWinTT [13], and GRM [22] introduce the transformer for feature interaction. Recently, a single-stream framework, exemplified by such as MixFormer [9], OSTrack [10], SimTrack [14], DropTrack [23], CTTrack [24], and ARTrack [25], has demonstrated great success in visual tracking. This framework combines feature extraction and fusion within a single backbone network. It is a simple yet effective approach that capitalizes on the capabilities of pre-trained image classification models. However, these methods were primarily designed for powerful GPUs, resulting in slow speeds on edge devices, which limits their practical applications. In this work, we also adopt this single-stream framework but with a focus on improving its efficiency. This emphasis on efficiency aims to make the framework more suitable for deployment on resource-constrained devices, enhancing its practical utility across various real-world applications.
@@ 2.2 Efficient Tracking Methods
Practical applications necessitate the use of efficient tracking systems capable of delivering exceptional performance at rapid speeds, especially when deployed on edge devices. While early techniques like ECO [26] and ATOM [27] were able to achieve real-time tracking on edge devices, their performance fell short in comparison to state-of-the-art trackers available today. Recent advancements have led to the development of more efficient tracking solutions. For instance, LightTrack [28] employs neural architecture search (NAS) to discover network configurations, resulting in a low computational overhead and relatively high performance. FEAR [29] introduces a family of efficient and accurate tracking methods through the utilization of a dual-template representation and a pixel-wise fusion block. Recently, HiT [30] presents a family of efficient transformer-based tracking models that use a bridge module and dual-image position encoding to bridge the gap between lightweight transformers and the tracking framework. MixFormerV2  [31] proposes a fully transformer-based tracking framework without any dense convolutional operations or complex score prediction modules, additionally, it employs a distillation-based model compression paradigm to enhance its efficiency. DyTrack [32] introduces a dynamic transformer framework that automatically configures reasoning paths based on the complexity of the input to make better use of computational resources. Despite significant advancements, these methods are highly tailored and have problems with universality. A widely applicable method that can be easily integrated into existing trackers is more desirable. In this work, we present a novel approach that is straightforward and can be seamlessly integrated into existing ViT-based methods.
@@ 2.3 Efficient Vision Transformers
ViT [15] revolutionized the field of computer vision with its impressive performance [33, 17, 16]. However, despite their powerful modeling capabilities, they have been facing challenges associated with speed limitations. Especially on resource-constrained edge devices, the demand for efficient ViTs intensifies. To expedite ViTs, various lightweight ViTs have been recently introduced, employing techniques like low-rank methods, model compression, and hybrid designs [34, 35, 36]. However, low-rank and quantization-based ViTs often sacrifice considerable accuracy for efficiency, while Hybrid ViTs with CNN-based components exhibit restricted input flexibility. In this work, we introduce a novel ViT pruning method based on local ranking to reduce the dimension of the latent representation of tokens to boost efficiency. On the other hand, recent endeavors in efficient ViTs based on conditional computation have explored adaptive inference methods for accelerated modeling. For instance, DynamicViT [37] introduces extra control gates to pause specific tokens using the Gumbel-softmax trick. DVT [38] recognizes that different images require varying numbers of tokens, and it automatically configures an appropriate number of tokens for each input image to improve computational efficiency. A-ViT [39] uses an Adaptive Computation Time (ACT)-like method to enhance efficiency, accuracy, and token importance allocation without requiring additional halting sub-networks. TPS [41] proposes a joint Token Pruning and Squeezing module for compressing vision transformers with higher efficiency. However, variable number of tokens in these methods incurs considerable time cost due to additional unstructured assess operations. Very recently, LGViT [40] accelerates the inference process of ViT through a dynamic early exiting mechanism, which is a more structured approach employing dynamic ViT blocks rather than varying the token number. Nevertheless, its training process is highly complex, and it cannot trim non-adjacent ViT blocks due to the constraints of the early exiting mechanism. In this study, we investigate adaptively bypassing ViT blocks as a structured and more effective method to enhance ViTs’ efficiency. As the tracking challenge for a specific target is highly context-dependent, adaptive bypassing of ViT blocks can notably decrease insignificant semantic features or relations at specific abstraction levels without significantly impacting the tracker’s performance.
3 Proposed Approach
In this section, we begin with a brief overview of our tracking framework, ABTrack, as illustrated in Figure 2. Next, we introduce the bypass decision module and the approach to pruning vision transformer models. Finally, we provide details on the prediction head and training objective.
@@ 3.1 Overview
Figure 2: (Left) Overview of the proposed ABTrack framework. It contains a single-stream backbone and a prediction head, in which the backbone consists of pruned ViT blocks and Bypass Decision Modules. (Right) The implementation details of the proposed Bypass Decision Module.
The proposed ABTrack is a single-stream tracking framework, consisting of an adaptive ViT-based backbone and a prediction head. The inputs to the framework are the target template and the search image. The extracted features from the backbone are fed into the prediction head to obtain tracking predictions. In the backbone network, we introduce a bypass decision module (BDM) for each ViT block (layer) except for the first $n_{enf}$ layers. Each BDM is trained to adaptively determines whether to bypass the associated block based on the computed bypassing probability. Furthermore, we conduct pruning on the ViT blocks using two introduced dimension reduction (DR) matrices. These matrices indicate the importance scores for each dimension of the linear layer of the Multi-Head Self-Attention (MSA) and Multi-Layer Perceptron (MLP) within each block, respectively. In our approach, the BDM maintains the innate transformation characteristics of each ViT layer to ensure consistent information processing throughout the network. However, ViT pruning modifies the BDM’s input domain, directly influencing its operational efficiency and performance. Therefore, we prioritize ViT pruning before integrating the BDMs, aiming to improve overall tracking accuracy and efficiency in visual tasks. Finally, with BDMs integrated, the pruned model undergoes fine-tuning within the tracking framework.
@@ 3.2 Bypass Decision Module (BDM)
The objective of the BDMs is to dynamically determine whether a transformer block should be bypassed based on the given template and search image. In consideration of efficiency, it is implemented by a linear layer followed by a nonlinear activation function. Its input is a bypass token, a specific summary token that is appended to the sequence of image tokens. And its output represents the probability of bypassing the associated transformer block. Specifically, take the $i_{th}$ ($i>n_{enf}$) layer for example, let the bypass token be denoted by $(t^{i}_{1},...,t^{i}_{d}):=\textbf{b}^{i}\in\mathbb{R}^{d}$, the linear layer by $\mathfrak{l}^{i}$, where $d$ is the token dimension, then formally the BDM at layer $i$ is defined by
$p_{i}=\sigma(\mathfrak{l}^{i}(\textbf{b}^{i})),$ (1)
@@
where $p_{i}\in[0,1]$ represents the probability of bypassing the $i$ transformer block, $\sigma(x)=1/(1+e^{-x})$ denotes the sigmoid function. If $p_{i}>\rho$, where $\rho\in[0,1]$ is the bypass probability threshold, the transformer block at layer $i$ is bypassed; otherwise, the output tokens from the $(i-1)_{th}$ layer is fed into the $i_{th}$ transformer block. Let $N$ denote the total number of transformer blocks of the given ViT. Theoretically, all $N$ blocks could be bypassed simultaneously, resulting in no correlation being computed between the template and the search image. To prevent such situations, the first $n_{enf}$ layers are enforced without bypassing. These low-level blocks are crucial because they provide foundational information that supports the construction of more complex features and representations. By prioritizing low-level features, the system ensures that the basic, yet critical, elements of the data are processed sufficiently, thus facilitating the subsequent development of higher-level abstractions. This practice also alleviates the computational burden caused by an excessive number of BDMs. Another extreme case occurs when few or no blocks are bypassed regardless of the input. This situation enhances the model’s representation ability, allowing it to capture more detailed and complex features. Consequently, this increases the model’s capacity to reduce training loss effectively. However, this can lead to higher computational costs and may not be efficient for real-time applications. Balancing between bypassing and retaining blocks is crucial to optimize both performance and efficiency. To this end, we propose an adaptive block sparsity loss, denoted as $\mathcal{L}_{spar}$. This loss function adaptively penalizes smaller average bypass probabilities across all layers considered in the model. The penalty is stricter when the tracking task is deemed easy and less strict when the task is challenging. This adaptive approach dynamically adjusts block bypassing, aiming to strike a balance between efficiency and accuracy tailored to the complexity of each tracking scenario. The block sparsity loss is formally defined by
@@
$\begin{split}&\mathcal{L}_{spar}=|\frac{1}{N-n_{enf}}\sum^{N}_{i=n_{enf}+1}p_{i}-\tau(\mathcal{L}_{iou};\overline{\mathcal{L}_{iou}})|,\\ &\tau(\mathcal{L}_{iou};\overline{\mathcal{L}_{iou}})=\text{clip}(\tau_{0}+\zeta(\mathcal{L}_{iou}-\overline{\mathcal{L}_{iou}}),0,1),\end{split}$ (2)
@@
where $\text{clip}(x,a,b)$ is a mathematical operation that limits the value of $x$ within a specified range defined by $a$ and $b$. $\tau\in[0,1]$ is a function of GIoU loss [42] (see Section 3.4) and is used along with $\rho$ to control the block sparsity. Note that a linear function of the deviation of GIoU loss from its average is used to quantify the difficulty of the tracking task. In our implementation, the average GIoU loss, i.e., $\overline{\mathcal{L}_{iou}}$, is estimated with a batch training samples. $\zeta>0$ and $\tau_{0}\in[0,1]$ are two constants used to model the linear relation between $\tau$ and $(\mathcal{L}_{iou}-\overline{\mathcal{L}_{iou}})$. We call $\tau$ the block sparsity function. Given $\rho$, generally the larger $\tau$ is, the more sparse the model will be.